determine the nature of the solutions of the equation
x^2+4=0
Please help me solve.

Answer by stanbon(75887)

determine the nature of the solutions of the equation
x^2+4=0
----
It's a quadratic with a = 1, b = 0, c = 4
----
The discriminant is b^2-4ac = 0^2 -4*1*4 = -16
---
Since the discriminant is negative, the quadratic
has two unequal complex solutions.
------------------------------------
What are they:
x^2+4 = 0
Factor:
(x+2i)(x-2i) = 0
x = -2i or x = +2i
=====================
Cheers,
Stan H.